Solve algebraically giving answers in terms of π:
2sinx=1, 0 ≤ x ≤ 6π
I don't understand the steps on how to solve it. Should I divide 2 with the radian for 1?

Answer by stanbon(75887)

Solve algebraically giving answers in terms of π:
2sinx=1, 0 ≤ x ≤ 6π
----------
sin(x) = 1/2,
---
Since x is limited to 0<= x <= 6*pi
---
x = (pi/6)+ 2n(pi) where n = 0,1,2
or
x = (5/6)pi + 2n(pi) where n = 0,1,2
----------------
Cheers,
Stan H.